A calorimeter of heat capacity 100J/K is at room temperature of 30’C.100g of water at 40’C of specific heat 4200J/Kg-K is poured into the calorimeter. What is the temperature of water in calorimeter?

Grade:12th pass

1 Answers

Rohan Rai
5 years ago
Let The Temperature of Water In Caloriemeter Is t.We know Heat lost by water = Heat Gained by CaloriemeterSo(0.1)×4200×(40-t) = 100(t-30)So t = 38.7°C